OpenStudy (anonymous):

There are 27 students in a chemistry class and 22 students in a physics class. Seven of these students take physics and chemistry. What is the ratio of the number of students taking only physics to those taking only chemistry? (A)4:3 (B)34:29 (C)7:6 (D)3:4 (E)22:27 I suck at ratios =_="

OpenStudy (blurbendy):

20/15 = 4/3 = 4:3 20/15 because you have 27 in chem and 2 in physics physics, but 7 are taking chem and physics, so 27 - 7= 20 and 22 -7 = 15

OpenStudy (anonymous):

Thanks.

OpenStudy (blurbendy):

yw